OpenStudy (anonymous):

Solve For x 2 ln 3x=5

OpenStudy (anonymous):

\[\ln(3x)=\frac{5}{2}\] \[e^{\frac{5}{2}}=3x\] \[x=\frac{1}{3}e^{\frac{5}{2}}\]